OpenStudy (anonymous):

help please! find Sn for the given geometric series.. a1=-10, a5=-12960, r=6 a. -15550 b.-32425 c.-11110 d. -2590

OpenStudy (btaylor):

of just the first 5 terms?

OpenStudy (amistre64):

use the formula \[a_1\frac{1-r^n}{1-r}\]

OpenStudy (anonymous):

how do i find n amistre64?

OpenStudy (amistre64):

n = number of terms; so there are 5 terms

OpenStudy (anonymous):

oo okay

OpenStudy (amistre64):

a1, a2, a3, a4, a5 ... n=5 :)

OpenStudy (anonymous):

hmm well i got -7781 with that formula...which is obviously wrong...what did i do?!

OpenStudy (amistre64):

did you keep track of your signs?

OpenStudy (anonymous):

well let me try again...

OpenStudy (anonymous):

..well im doing something wrong...i plugged in -10=1-6^5/1-6...i didn't include -12960..which is probably where im going wrong

OpenStudy (amistre64):

the equal sign is not a good thing

OpenStudy (amistre64):

-10 * (1-6^5) ------------ (1-6)

OpenStudy (anonymous):

ooo okay multiply not equals..makes sense! thank you!

OpenStudy (amistre64):

1-6 = -5 so -10/-5 = 2 for simplicities sake

OpenStudy (amistre64):

2(1-6^5)

OpenStudy (anonymous):

ok i see now...thanks!

OpenStudy (amistre64):

yw, and good luck ;)

OpenStudy (anonymous):

so wats the answer

OpenStudy (amistre64):

@e.mccormick hmmm, this looks familiar :)

OpenStudy (e.mccormick):

@dzenan The answer is to look at that formula, \(a_1\frac{1-r^n}{1-r}\), and learn how to use it. If you have trouble, ask your own question about how to use the formula and then you can get the answer!